Problem 22 Consider the following forecast results. Caloulate MFE, MAD, and MAPE, using the data for the months January through June. Does the forecast model under- or overforecast? Actual Forecast Month Demand January February March 968 1,091 1,089 1,063 1,054 898 1,029 1,113 June 1,013 1.100 The MFE slag (Enter your response rounded to one decimal place and include The MAD is(Enter your response rounded to one decimal place,) a minus sign if necessary)
Answer #1

Please provide rating..

MFE =sum of (Actual-Forecast)/Number of month
MAD = sum of|Actual-forecast|/Number of month
MAPE = sum of(|Actual-Forecast|/Actuals demand)/number of month
i ii iii=i-ii iv=|i-ii| v=iv/i
Month Actual demand Forecast MFE MAD MAPE
January 1018 1063      (45.00)                         45.00 4.42%
February 968 1054      (86.00)                         86.00 8.88%
March 963 898       65.00                         65.00 6.75%
April 1091 1029       62.00                         62.00 5.68%
May 1089 1113      (24.00)                         24.00 2.20%
June 1013 1100      (87.00)                         87.00 8.59%
Total (115.00)                       369.00 36.53%
Average      (19.17)                         61.50 6.09%
MFE MAD MAPE
MFE          (19.17)
MAD            61.50
MAPE 6.09%
The forecast model is a overforecast as MFE is negative